A number, 1/2 of that number and 1/3 of that number added. The result is 22.what is the original number.
Found 2 solutions by solver91311, checkley77:
Answer by solver91311(24713)





Solve for .

John

My calculator said it, I believe it, that settles it
The Out Campaign: Scarlet Letter of Atheism


Answer by checkley77(12844)

N+N/2+N/3=22
(6N+3N+2N)/6=22
11N=132N=132/11
N=12 ANS.
PROOF:
12+12/2+12/3=22
12+6+4=22
22=22